# Bier spheres and Toric Topology

### v1

@@ Abstract.
We compute the real and complex Buchstaber numbers of an arbitrary Bier sphere. In dimension two, we identify all the 13 different combinatorial types of Bier spheres and show that 11 of them are nerve complexes of nestohedra, while the remaining two are merely nerve complexes of generalized permutohedra: one of them is a non-flag polytope and the other one is a 2-truncated cube. As an application of our results, we construct a regular normal fan for each of those 13 Delzant polytopes, compute cohomology rings of the corresponding nonsingular projective toric varieties and discuss orientability of the corresponding small covers.
@@ 1. Introduction
Some of the known applications of Bier spheres in topological combinatorics are related to the fact that they have become a useful tool there thanks to the connection of their theory with the Van Kampen–Flores theorem, see [17]. In addition, it was found in [12, 13, 14] that any Bier sphere admits a starshaped realization. Therefore, each Bier sphere corresponds to at least one complete toric variety (possibly a singular one).
@@
In Section 2 we recall all the necessary definitions related to simplicial complexes and simple polytopes. Then we show that there exist only 13 different combinatorial types of Bier spheres in dimension two. Since any 2-dimensional triangulated sphere is polytopal, we are able to sketch them as boundaries of the 3-dimensional simplicial polytopes and find geometric realizations of their dual simple polytopes as generalized truncation polytopes and nestohedra, whenever possible. In particular, we show that all the 13 simple polytopes arising in this way have Delzant realizations, all but two of them are nestohedra, and only 4 of them are flag polytopes (Theorem 2.18).
@@
In Section 4 we consider some examples of our main results application. We briefly recall the notions of a quasitoric manifold and a small cover and construct the canonical regular normal fans for each of the 11 nestohedra provided by our Bier spheres classification using the result of [11]. Then we compute cohomology rings of the resulting 13 nonsingular projective toric varieties (Example 4.5) and identify orientable manifolds among the resulting 13 small covers (Example 4.6).
@@ 2.1. Preliminaries
The latter notion was introduced in [2], where it was also proved that $\mathrm{Bier}(\mathcal{K})$ is an $(m-2)$-dimensional PL-sphere with a number of vertices varying between $m$ and $2m$, for any simplicial complex $\mathcal{K}\neq\Delta_{[m]}$ on $[m]$. This means that there exists a subdivision of the $(m-2)$-dimensional triangulated sphere $\mathrm{Bier}(\mathcal{K})$ combinatorially equivalent to a subdivision of $\partial\Delta^{m-1}$, the boundary of the $(m-1)$-dimensional simplex.
@@ 2.2. Combinatorial types of $\mathrm{Bier}_{4}(\mathcal{K})$
Let us describe the minimal non-faces of $\mathcal{K}*_{\Delta}\widehat{\mathcal{K}}$. These are $\sigma\uplus\tau$, which is either $\sigma\uplus\tau\notin\mathcal{K}*\widehat{\mathcal{K}}$, or $\sigma\uplus\tau\in\mathcal{K}*\widehat{\mathcal{K}}$ and $\sigma\cap\tau\neq\varnothing$.
If $\sigma\uplus\tau\in\mathcal{K}*\widehat{\mathcal{K}}$ and $\sigma\cap\tau\neq\varnothing$, then the minimality of this non-face means that any proper face of $\sigma$ does not intersect any (not necessarily proper) face of $\tau$ and vice versa. This is possible only if $\sigma=\tau=\{i\}$, where $\{i\}$ is a vertex belonging to $\mathcal{K}$ and $\widehat{\mathcal{K}}$. Thus, we get $\{x_{i},y_{i}\}=x_{i}y_{i}\in\mathrm{MF}(\mathrm{Bier}(\mathcal{K}))$. Note that if $\{i\}$ is a ghost vertex for $\mathcal{K}$ or $\widehat{\mathcal{K}}$, then $x_{i}y_{i}\notin\mathrm{MF}(\mathrm{Bier}(\mathcal{K}))$.
If $\sigma\uplus\tau\notin\mathcal{K}*\widehat{\mathcal{K}}$, then, given that at least one of them must be a simplex, we get $\sigma\notin\mathcal{K}$ and $\tau\in\widehat{\mathcal{K}}$, or $\sigma\in\mathcal{K}$ and $\tau\notin\widehat{\mathcal{K}}$. In the first case, the minimality means $\sigma\in\mathrm{MF}(\mathcal{K})$, $\tau=\varnothing$, and any proper face of $\sigma$ does not intersect $\tau$ automatically. In the second case, the minimality similarly means $\tau\in\mathrm{MF}(\widehat{\mathcal{K}})$ and $\sigma=\varnothing$. ∎
@@
Let $\mathcal{K}$ be an arbitrary simplicial complex on $V$. Then the number of ghost vertices of $\mathrm{Bier}(\mathcal{K})$ is equal to $|V|-f_{0}(\mathcal{K})+f_{f_{0}(\mathcal{K})-2}(\mathcal{K})$.
@@
Indeed, the ghost vertices of the Bier sphere $\mathrm{Bier}(\mathcal{K})$ are exactly the singleton elements of $\mathrm{MF}(\mathrm{Bier}(\mathcal{K}))$. Those, in turn, are the singleton elements of either $\mathrm{MF}(\mathcal{K})$ or $\mathrm{MF}(\widehat{\mathcal{K}})$, by the previous result. By definition, simplices of $\widehat{\mathcal{K}}$ are complements to non-simplices of $\mathcal{K}$. Thus, ghost vertices of $\widehat{\mathcal{K}}$ are exactly the complements to the faces of $\mathcal{K}$ of cardinality $f_{0}(\mathcal{K})-1$. ∎
@@
Table 1. Simplicial complexes on four vertices and their Alexander duals. The painted dots mean vertices and the unpainted ones mean ghost vertices
@@ 2.3. Affine realizations and generalized truncation polytopes
Furthermore, $\mathcal{P}_{7}$ is $vc^{1}(\Delta^{1}\times\Delta^{1}\times\Delta^{1})$, $\mathcal{P}_{8}$ is $2$-truncated cube (with one $2$-truncation performed), $\mathcal{P}_{9}$ is $vc^{3}(\Delta^{3})$, $\mathcal{P}_{10}=\Delta^{1}\times\Delta^{1}\times\Delta^{1}$, $\mathcal{P}_{11}$ is a $2$-truncated triangular prism (first the vertex of $\Delta^{3}$ is cut off, then the edge from the base), $\mathcal{P}_{12}$ is $vc^{1}(\Delta^{3})$, $\mathcal{P}_{13}=\Delta^{3}$.
@@ 2.4. Delzant realizations and nestohedra
A Delzant polytope is a lattice polytope such that its normal fan is regular. Two lattice polytopes are equivalent if their normal fans are isomorphic (i.e. they can be mapped to each other by the action of $GL_{n}(\mathbb{Z})$). A Delzant realization of a simple polytope $P$ is an arbitrary Delzant polytope which is combinatorially equivalent to $P$.
It was shown in [9] that if a 3-polytope has a Delzant realization, then it has at least one triangular or quadrangular facet. Two relevant and particularly important classes of Delzant polytopes are given by 2-truncated cubes and nestohedra.
@@
Proposition 2.11 ([20, 5]).
Let $\mathcal{B}$ be a building set on $[n+1]$. We have
$P_{\mathcal{B}}=\Big{\{}\textbf{x}\in\mathbb{R}^{n+1}\,\big{|}\,\sum^{n+1}_{i=1}x_{i}=|\mathcal{B}|,\;\sum_{i\in S}x_{i}\geqslant|\mathcal{B}|_{S}|\;\text{for every }S\in\mathcal{B}\Big{\}},$
where $\mathcal{B}|_{S}=\{S^{\prime}\in\mathcal{B}\emph{\text{ | }}S^{\prime}\subset S\}$. If $\mathcal{B}$ is connected, then this representation is irredundant, that is, every hyperplane
$H_{S}=\Big{\{}\textbf{x}\in\mathbb{R}^{n+1}\,\big{|}\,\sum_{i\in S}x_{i}=|\mathcal{B}|_{S}|\Big{\}}$
with $S\neq[n+1]$ defines a facet $F_{S}$ of $P_{\mathcal{B}}$ (so that the number of facets of $P_{\mathcal{B}}$ is $|\mathcal{B}|-1$).
@@
It follows from the last construction that any nestohedron can be obtained by face truncations of a simplex. Hence, every nestohedron is a Delzant polytope. Besides, two combinatorially equivalent nestohedra are equivalent as lattice polytopes. So, every simple polytope $P$ is either combinatorially equivalent to some nestohedron (in this situation the nestohedron is well-defined and appears to be the “canonical” Delzant realization of $P$), or does not admit a realization as a nestohedron (then a Delzant realization might not exist).
@@
Now, we are going to show that all the 3-dimensional Bier polytopes are nestohedra, except for two, which are merely a generalized permutohedra.
@@
It is well-known that every $n$-dimensional nestohedron can be obtained from the $n$-simplex [^1] In the flag case every $n$-dimensional nestohedron is an $n$-dimensional 2-truncated cube, see [6, Theorem 6.6] by a sequence of face truncations, see [5, Theorem 1.5.18].
Firstly, note that we cannot get the $\mathcal{P}_{6}$ by face truncations of the cube $I^{3}$ with a nestohedron realization given by the building set $\mathcal{B}_{I^{3}}=\Big{\{}\{1\},\{2\},\{3\},\{4\},\{1,2\},\{1,2,3\},\{1,2,3,4\}\Big{\}}$, since it is not allowed to cut any vertex of the face corresponding to $\{1,2,3\}\in\mathcal{B}_{I^{3}}$. We will get into this situation if we first cut off the vertex, and then the edge incident with it (see Figure 1).
Figure 1. $\mathcal{B}_{I^{3}}=\Big{\{}\{1\},\{2\},\{3\},\{4\},\{1,2\},\{1,2,3\},\{1,2,3,4\}\Big{\}}$
@@
Secondly, note that cutting off the vertex of a simple $3$-polytope adds $1$ triangular face, $3$ edges and $2$ vertices, and slicing an edge adds $1$ quadrangular face, $3$ edges and $2$ vertices. The $f$-vector of $\mathcal{P}_{6}$ is $(12,18,8)$ and the $f$-vector of $\Delta^{3}$ is $(4,6,4)$. Thus, it is necessary to do exactly $\frac{12-4}{2}=4$ cuts to get $\mathcal{P}_{6}$ from $\Delta^{3}$. Observe that the $\mathcal{P}_{6}$ does not have a quadrangular face. So the last step should be the vertex truncation.
If the first step is to cut off a vertex of the simplex, we get a triangular prism. As we have already noticed above, we must not cut off the side edges of this prism. In this case, we cannot cut off a vertex of the newly formed face. Therefore, in the second step we cut off a vertex or an edge of another base.
(a) Hexagonal face after $3$ vertex truncations.
(b) Edge truncation of the triangular prism.
Figure 2.
If we cut off a vertex, we get two triangular faces, each of which cannot have vertices cut off. If we cut off a vertex again, a hexagonal face will be formed (see Figure 2 (a)). Since we cannot cut edges of the newly formed triangular faces, we need to study a truncation of vertices or edges of the edge-truncated triangular prism (see Figure 2 (b)). Note that in this case we can only truncate one vertex. If we do that, a hexagonal face will be formed again.
If the first step is to cut off the edge of the simplex, we arrive at Figure 3. We have already analyzed the cases of truncation of its vertices in the second step (Figure 2 (b)). Recall that the last step should be to truncate a vertex. Since we cannot cut off the vertices or edges adjacent to the vertices of the simplex, one of the vertices of the base can be considered truncated, which leads to the cases already considered, which completes the proof.
Figure 3. Edge truncation of $\Delta^{3}$
@@
Still, $\mathcal{P}_{6}$ has a realization with a regular normal fan, as the next result shows.
@@
Theorem 2.16.
$\mathcal{P}_{4}$ is not a nestohedron.
Proof.
Assuming the opposite, let $\mathcal{B}\subset 2^{[4]}$ be a connected building set for the flag polytope $\mathcal{P}_{4}$. Note that there are only two ways to obtain the cube $I^{3}$ by cutting off the faces of a simplex in the three-dimensional case: by cutting off the vertex and then the edge (see Figure 1) or by cutting off the edge of the edge-truncated simplex (see Figure 3). Therefore, without loss of generality, we can assume that the building set of the $3$-cube has the form
Every $n$-dimensional flag nestohedron is an $n$-dimensional 2-truncated cube, see [6, Theorem 6.6]; that is $\mathcal{B}_{I^{3}}\subset\mathcal{B}$ and all truncations are edge-truncations with respect to $\mathcal{B}_{I^{3}}$. Recall that $\mathcal{P}_{4}$ is formed by cutting off two adjacent edges of the $3$-cube. Since we must not cut off edges adjacent to the face corresponding to $\{1,2,3\}$, it only remains to cut off the faces corresponding to $\{1,3\}$ and $\{3,4\}$. But in this case we will also have to add $\{1,3,4\}$ and we obtain
=$\mathcal{B}=\Big{\{}\{1\},\{2\},\{3\},\{4\},\{1,2\},\{1,2,3\},\{3,4\},\{1,3\},\{1,3,4\},\{1,2,3,4\}\Big{\}}.$
Therefore, we have added an “extra edge”, a contradiction. ∎
Proposition 2.17.
$\mathcal{P}_{4}$ admits a Delzant realization in $\mathbb{R}^{3}$.
Proof.
Let $(a,b,c)\in\mathbb{R}^{3}$ be the canonical coordinates. Define a cube $C$ using inequalities
$0\leqslant a\leqslant 4,\quad 0\leqslant b\leqslant 4,\quad 0\leqslant c\leqslant 4,$
and cut it by the two hyperplanes
$a+b-6=0,\quad a+c-7=0.$
We have obtained some realization of $\mathcal{P}_{4}$ in $3$-space. Now we will show that this is a Delzant realization. Consider the normal vectors to the facets as columns of the following integer matrix
$\begin{pmatrix}1&1&0&-1&0&0&0&1\\ 0&0&0&0&0&-1&1&1\\ 0&1&1&0&-1&0&0&0\end{pmatrix}\in\mathrm{Mat}_{3\times 8}(\mathbb{Z}).$
It is easy to see that the columns of this matrix serve as generators of a complete regular fan, and therefore, we obtained a Delzant realization of the corresponding Bier polytope. ∎
@@
There are exactly $13$ pairwisely non-isomorphic two-dimensional Bier spheres $\mathcal{S}_{i}$. As for the corresponding Bier polytopes $\mathcal{P}_{i}$:
@@
    For each $i=1,\ldots,3,\not$ /$4$,5,,5$\not$, 5 , /$6$,7,,13,7$\ldots$,13, 7 , … , 13 the polytope $\mathcal{P}_{i}$ is a nestohedron;
@@
  • $\mathcal{B}_{4}\quad$ does not exist by Theorem 2.16;
@@
By Propositions 2.15 and 2.17, polytopes $\mathcal{P}_{4}$ and $\mathcal{P}_{6}$ are not nestohedra, but admit Delzant realisations in $\mathbb{R}^{3}$. Q.E.D. ∎

### v2

@@ Abstract.
We compute the real and complex Buchstaber numbers of an arbitrary Bier sphere. In dimension two, we identify all the 13 different combinatorial types of Bier spheres and show that 12 of them are nerve complexes of nestohedra, while the remaining one is a nerve complex of a generalized permutohedron. As an application of our results, we construct a regular normal fan for each of those 13 Delzant polytopes, compute the cohomology rings of the corresponding nonsingular projective toric varieties, and examine the orientability of the corresponding small covers.
@@ 1. Introduction
Some of the known applications of Bier spheres in topological combinatorics are related to the fact that they have become a useful tool there thanks to their connection of their theory with the Van Kampen–Flores theorem, see [17]. In addition, it was found in [13] that any Bier sphere admits a starshaped realization. Therefore, each Bier sphere corresponds to at least one complete toric variety (possibly a singular one).
@@
In Section 2 we recall all the necessary definitions related to simplicial complexes and simple polytopes. Then we show that there exist only 13 different combinatorial types of Bier spheres in dimension two. Since any 2-dimensional triangulated sphere is polytopal, we are able to sketch them as boundaries of the 3-dimensional simplicial polytopes and find geometric realizations of their dual simple polytopes as generalized truncation polytopes and nestohedra, whenever possible. In particular, we show that all the 13 simple polytopes arising in this way have Delzant realizations, all but one of them are nestohedra, and only 4 of them are flag polytopes (Theorem 2.16).
@@
In Section 4 we consider some examples of our main results application. We briefly recall the notions of a quasitoric manifold and a small cover and construct the canonical regular normal fans for each of the 12 nestohedra provided by our Bier spheres classification using the result of [11]. Then we compute cohomology rings of the resulting 13 nonsingular projective toric varieties (Example 4.5) and identify orientable manifolds among the resulting 13 small covers (Example 4.6).
@@ 2.1. Preliminaries
The latter construction was introduced in the unpublished note  [2], where it was also proved that $\mathrm{Bier}(\mathcal{K})$ is an $(m-2)$-dimensional PL-sphere with a number of vertices varying between $m$ and $2m$, for any simplicial complex $\mathcal{K}\neq\Delta_{[m]}$ on $[m]$. This means that there exists a subdivision of the $(m-2)$-dimensional triangulated sphere $\mathrm{Bier}(\mathcal{K})$ combinatorially equivalent to a subdivision of $\partial\Delta^{m-1}$, the boundary of the $(m-1)$-dimensional simplex.
@@ 2.2. Combinatorial types of $\mathrm{Bier}_{4}(\mathcal{K})$
Let us describe the minimal non-faces of $\mathcal{K}*_{\Delta}\widehat{\mathcal{K}}$. These are $\sigma\uplus\tau$ such that either $\sigma\uplus\tau\notin\mathcal{K}*\widehat{\mathcal{K}}$, or $\sigma\uplus\tau\in\mathcal{K}*\widehat{\mathcal{K}}$ and $\sigma\cap\tau\neq\varnothing$.
If $\sigma\uplus\tau\in\mathcal{K}*\widehat{\mathcal{K}}$ and $\sigma\cap\tau\neq\varnothing$, then the minimality of this non-face means that any proper face of $\sigma$ does not intersect any (not necessarily proper) face of $\tau$ and vice versa. This is possible if and only if $\sigma=\tau=\{i\}$ for a vertex $\{i\}$ that belongs to $\mathcal{K}$ and $\widehat{\mathcal{K}}$. Thus, we get $\{x_{i},y_{i}\}=x_{i}y_{i}\in\mathrm{MF}(\mathrm{Bier}(\mathcal{K}))$. Note that if $\{i\}$ is a ghost vertex for $\mathcal{K}$ or $\widehat{\mathcal{K}}$, then $x_{i}y_{i}\notin\mathrm{MF}(\mathrm{Bier}(\mathcal{K}))$.
If $\sigma\uplus\tau\notin\mathcal{K}*\widehat{\mathcal{K}}$, then we get either $\sigma\notin\mathcal{K}$ and $\tau\in\widehat{\mathcal{K}}$, or $\sigma\in\mathcal{K}$ and $\tau\notin\widehat{\mathcal{K}}$. In the former case, the minimality means $\sigma\in\mathrm{MF}(\mathcal{K})$, $\tau=\varnothing$, and any proper face of $\sigma$ does not intersect $\tau$ automatically. In the latter case, the minimality similarly means that $\tau\in\mathrm{MF}(\widehat{\mathcal{K}})$ and $\sigma=\varnothing$. ∎
@@
Let $\mathcal{K}$ be an arbitrary simplicial complex on $V$. Then the number of ghost vertices of $\mathrm{Bier}(\mathcal{K})$ is equal to $|V|-f_{0}(\mathcal{K})+f_{|V|-2}(\mathcal{K})$.
@@
Indeed, the ghost vertices of the Bier sphere $\mathrm{Bier}(\mathcal{K})$ are exactly the singleton elements of $\mathrm{MF}(\mathrm{Bier}(\mathcal{K}))$. Those, in turn, are the singleton elements of either $\mathrm{MF}(\mathcal{K})$ or $\mathrm{MF}(\widehat{\mathcal{K}})$, by the previous result. By definition, the simplices of $\widehat{\mathcal{K}}$ are the complements of non-simplices of $\mathcal{K}$. Thus, the ghost vertices of $\widehat{\mathcal{K}}$ are exactly the complements of the faces of $\mathcal{K}$ of cardinality $|V|-1$. ∎
@@
Table 1. Simplicial complexes on four vertices and their Alexander duals. The solid dots mean vertices, and the hollow ones, ghost vertices.
@@ 2.3. Affine realizations and generalized truncation polytopes
Furthermore, $\mathcal{P}_{7}$ is $vc^{1}(\Delta^{1}\times\Delta^{1}\times\Delta^{1})$, $\mathcal{P}_{8}$ is $2$-truncated cube (with one $2$-truncation performed), $\mathcal{P}_{9}$ is $vc^{3}(\Delta^{3})$, $\mathcal{P}_{10}=\Delta^{1}\times\Delta^{1}\times\Delta^{1}$, $\mathcal{P}_{11}$ is $vc^{2}(\Delta^{3})$, $\mathcal{P}_{12}$ is $vc^{1}(\Delta^{3})$, $\mathcal{P}_{13}=\Delta^{3}$.
@@ 2.4. Delzant realizations and nestohedra
A Delzant polytope is a lattice polytope such that its normal fan is regular; i.e., the primitive normal vectors to facets meeting at a vertex form a lattice basis for each vertex of the lattice polytope. A Delzant realization of a simple polytope $P$ is an arbitrary Delzant polytope which is combinatorially equivalent to $P$.
It was shown in [9] that if a 3-polytope has a Delzant realization, then it has at least one triangular or quadrangular facet. Two classes of Delzant polytopes are extremely important for our classification: nestohedra and 2-truncated cubes.
@@
Proposition 2.11 ([20]).
Let $\mathcal{B}$ be a connected building set on $[n+1]$. Then nestohedron $P_{\mathcal{B}}$ is an $n$-dimensional simple polytope equal to the intersection of the hyperplane $H=\{(x_{1},\dots,x_{n+1})\in\mathbb{R}^{n+1}\emph{\text{ | }}\sum^{n+1}_{i=1}x_{i}=|\mathcal{B}|\}$ with the half space $H_{S}=\{(x_{1},\dots,x_{n+1})\in\mathbb{R}^{n+1}\emph{\text{ | }}\sum_{i\in S}x_{i}\geqslant|\mathcal{B}|_{S}|\}$, where $\mathcal{B}|_{S}=\{S^{\prime}\in\mathcal{B}\emph{\text{ | }}S^{\prime}\subset S\}$. Moreover, there is a one to one correspondence between facets of $P_{\mathcal{B}}$ and elements of \$\mathcal{B}\backslash\{[n+1]\}$, such that $S\in\mathcal{B}$ corresponds to the facet $F_{S}=P_{\mathcal{B}}\cap\partial H_{S}$.
@@
It follows from the last construction that any nestohedron can be obtained by face truncations of a simplex. Hence, every nestohedron is a Delzant polytope. So, every simple polytope $P$ is either combinatorially equivalent to some nestohedron (in this situation the nestohedron appears to be the “canonical” Delzant realization of $P$), or does not admit a realization as a nestohedron (then a Delzant realization might not exist).
@@
Now, we are going to show that all the 3-dimensional Bier polytopes are nestohedra, except for one, which is merely a generalized permutohedron.
@@
Suppose that $\mathcal{P}_{6}$ has a combinatorial type of nestohedron with a building set $\mathcal{B}$. Let
$\mathcal{B}_{\Delta^{3}}=\Big{\{}\{1\},\{2\},\{3\},\{4\},\{1,2,3,4\}\Big{\}}$
@@
be a building set of $\Delta^{3}$. It is well-known that every $n$-dimensional nestohedron can be obtained from the $n$-simplex [^1] In the flag case every $n$-dimensional nestohedron is an $n$-dimensional 2-truncated cube, see [6, Theorem 6.6] by a sequence of face truncations, see [5, Theorem 1.5.18]. In other words, $\mathcal{B}_{\Delta^{3}}\subset\mathcal{B}$.
@@
Note that cutting off the vertex of a simple $3$-polytope adds $1$ triangular face, $3$ edges and $2$ vertices, and slicing an edge adds $1$ quadrangular face, $3$ edges, and $2$ vertices. The $f$-vector of $\mathcal{P}_{6}$ is $(12,18,8)$ and the $f$-vector of $\Delta^{3}$ is $(4,6,4)$. Thus, it is necessary to do exactly $\frac{12-4}{2}=4$ cuts to get $\mathcal{P}_{6}$ from $\Delta^{3}$. Observe that $\mathcal{P}_{6}$ does not have a quadrangular face. So the last step should be the vertex truncation.
Case 1. $\mathcal{B}$ contains a building set $\mathcal{B}_{I^{3}}$ of the $3$-cube $I^{3}$. Without loss of generality, let
Figure 1. $\mathcal{B}_{I^{3}}=\Big{\{}\{1\},\{2\},\{3\},\{4\},\{1,2\},\{1,2,3\},\{1,2,3,4\}\Big{\}}$.
It follows from Lemma 2.12 that $\mathcal{P}_{6}$ can be obtained from $I^{3}$ by two truncations corresponding to some decompositions. The last step is to truncate some vertex. This means that \$\mathcal{B}\backslash\mathcal{B}_{I^{3}}=\big{\{}S^{1},S^{2}\big{\}}$ and $S^{2}=S_{1}^{2}\sqcup S^{2}_{2}\sqcup S^{2}_{3}$ has three disjoint elements of $\mathcal{B}_{I^{3}}$. Due to the minimality of decompositions, $\big{\{}S^{2}_{1},S^{2}_{2},S^{2}_{3}\big{\}}$ is
$\text{either}\quad\Big{\{}\{1\},\{3\},\{4\}\Big{\}}\quad\text{or}\quad\Big{\{}\{2\},\{3\},\{4\}\Big{\}}.$
If the decomposition of $S^{1}$ consists of only two elements, then there will be only one triangular face in $\mathcal{P}_{6}$. Therefore $\big{\{}S^{1},S^{2}\big{\}}=\Big{\{}\{1,3,4\},\{2,3,4\}\Big{\}}$, and we arrive to $\mathcal{P}_{2}$, which is not combinatorially equivalent to $\mathcal{P}_{6}$ by the classification of the Bier polytopes.
Case 2. $\mathcal{B}$ does not contain a building set of the $3$-cube $I^{3}$. Again, by Lemma 2.12, \$\mathcal{B}\backslash\mathcal{B}_{\Delta^{3}}=\big{\{}S^{1},S^{2},S^{3},S^{4}\big{\}}$ and $S^{4}$ corresponds to vertex truncation. Hence $S^{4}$ is decomposed as $S^{4}_{1}\sqcup S^{4}_{2}\sqcup S^{4}_{3}$, where $S^{4}_{i}\in\mathcal{B}_{\Delta^{3}}$, $i=1,2,3$. Observe that all $S^{4}_{i}$ are singletons. Therefore, if we add $S^{4}$ to $\mathcal{B}_{\Delta^{3}}$, we obtain a building set corresponding to the triangular prism $\Delta^{2}\times[0,1]$. Without loss of generality, let
$\mathcal{B}_{\Delta^{2}\times[0,1]}=\Big{\{}\{1\},\{2\},\{3\},\{4\},\{1,2,3\},\{1,2,3,4\}\Big{\}}.$
Since $\mathcal{P}_{6}$ has two triangular faces, we have to perform a vertex truncation once more. We arrive to $\mathcal{P}_{11}$. Due to symmetry, without loss of generality, assume that
$\mathcal{B}_{\mathcal{P}_{11}}=\Big{\{}\{1\},\{2\},\{3\},\{4\},\{1,2,3\},\{2,3,4\},\{1,2,3,4\}\Big{\}}.$
If we cut off a vertex again, a hexagonal face will be formed (Figure 2). By assumption, we can’t add $\{1,2\},\{1,3\},\{1,4\},\{2,3\},\{2,4\},\{3,4\}$ to $\mathcal{B}_{\mathcal{P}_{11}}$, a contradiction. Q.E.D.
Figure 2. Hexagonal face after $3$ vertex truncations.
@@
However, $\mathcal{P}_{6}$ has a realization with a regular normal fan, as the following result shows.
@@
There are exactly $13$ pairwisely non-isomorphic two-dimensional Bier spheres $\mathcal{S}_{i}$. The corresponding Bier polytopes $\mathcal{P}_{i}$ have the following properties:
@@
    For each $i=1,\ldots,5,\not$ /$6$,7,,13,7$\ldots$,13, 7 , … , 13 the polytope $\mathcal{P}_{i}$ is a nestohedron;
@@
  • $\mathcal{B}_{4}=\Big{\{}\{1\},\{2\},\{3\},\{4\},\{1,2\},\{1,2,3\},\{1,2,3,4\},\{3,4\},\{1,3,4\}\Big{\}}$;
@@
($\mathcal{B}_{6}$ does not exist by Theorem 2.14).
By Proposition 2.15, the polytope $\mathcal{P}_{6}$ is not a nestohedron but admits a Delzant realization in $\mathbb{R}^{3}$. ∎